OpenStudy (anonymous):

Find the area bounded by the curves x = 5 - y^2 and x - 2y = -5.

OpenStudy (shamil98):

first, find out where two curves intersect so set them equal to each other and solve for x.

OpenStudy (shamil98):

or you could solve for y and then find out the x value by plugging in.

OpenStudy (anonymous):

okay I graphed it. I found where the curves intersect

OpenStudy (anonymous):

(1,-2) and (5,0)

OpenStudy (anonymous):

k now i plug that into an integral

OpenStudy (shamil98):

Yeah. Remember area = top curve - bottom curve your bounds are from x = 1 to x = 5

OpenStudy (anonymous):

(12)-(12-8y)

OpenStudy (anonymous):

does that look right?

OpenStudy (shamil98):

nope

OpenStudy (anonymous):

k let me try again

OpenStudy (shamil98):

x = 5-y^2 x = 2y - 5 2y - 5 is the top curve 5-y^2 is the bottom curve change it to a function of y and then integrate it with respect to x

OpenStudy (anonymous):

so I have to get y on one side of the equation. and the x's on the other and then integrate

OpenStudy (shamil98):

yeah just change it

OpenStudy (shamil98):

5 - x = y^2 sqrt(5-x) = y x + 5 = 2y (x+5)/2 = y so the integral would be: \[\int\limits_{1}^{5} (\frac{ x+5 }{ 2 } - \sqrt{5-x}) dx\]

OpenStudy (anonymous):

I keep getting 10.66 but that is not the right answer.

OpenStudy (shamil98):

hm, maybe integrate it with respect to y?

OpenStudy (anonymous):

okay I will try that

OpenStudy (shamil98):

the bounds will be different,

OpenStudy (anonymous):

What will I change the bounds to?

OpenStudy (anonymous):

will it be -2 to 0

OpenStudy (anonymous):

the y coordinates

OpenStudy (shamil98):

yeah

OpenStudy (shamil98):

2y - 5 - (5-y^2) dy integrate that from -2 to 0